OpenStudy (youngstudier):

Simplify. 3 – [–(–1)] A. 4 B. –2 C. 2 D. –4 I think C.

OpenStudy (youngstudier):

@NathaliaMA13

OpenStudy (nathaliama13):

D. :)

OpenStudy (nathaliama13):

Each time you see an equation like this (for example) \[-6 - (-4)\] you add the equations :) instead of ACTUALLY subtracting.

OpenStudy (youngstudier):

but it's a positive 3...

OpenStudy (nathaliama13):

I know lmao. you still do the same thing when the number is either positive or negative. when you see double signs, you add. 3 -(-4) = 3 + 4 = 7 -3-(-4) = −3 + 4 = 1

OpenStudy (youngstudier):

Ok, thanks...

OpenStudy (nathaliama13):

I went over what I put as the answer..haha. D is def not the answer. it's C. my bad. really sorry.